A chemist must prepare 700.mL of 2.00M aqueous sodium thiosulfate Na2S2O3 working solution. He'll do this by pouring out some 3.15mol/L aqueous sodium thiosulfate stock solution into a graduated cylinder and diluting it with distilled water. Calculate the volume in mL of the sodium thiosulfate stock solution that the chemist should pour out. Round your answer to 3 significant digits.
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 3.15 M

M2 = 2 M

V2 = 700 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (2*700)/3.15

V1 = 444 mL

Answer: 444 mL